Circular No. 8468

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005K IN NGC 2923 Further to IAUC 8467, O. Trondal and M. Schwartz report the discov-

ery of an apparent supernova (mag ∼ 17.8) on unfiltered Tenagra II 0.81-m telescope images taken on Jan. 15.41 and 17.20 UT. SN 2005K is located at α = 9^{h}36^{m}04^{s}.31, δ = +16^{o}45′46′′.8 (equinox 2000.0), which is 6′′.7 east and 9′′.5 north of the nucleus of NGC 2923. Nothing is visible at this position on an image taken on 2004 Sept. 12.55 (limiting mag ∼ 19.0).

SUPERNOVAE 2004gx, 2004gz, 2005C, AND 2005G M. Ganeshalingam, F. J. D. Serduke, and A. V. Filippenko, University

of California, Berkeley, report that inspection of CCD spectra (range 330– 1060 nm), obtained on Jan. 16 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2004gz (IAUC 8460) is of type Ia, a few weeks past maximum brightness. SN 2005C (IAUC 8463) is probably of type Ib, several weeks past maximum; a set of He I absorption lines appears to be present. SN 2005G (IAUC 8465) is probably of type Ia near maximum, but the spectrum is somewhat peculiar (e.g., unusually narrow Si II 635.5-nm absorption; the two S II absorption lines near 550 nm are blended together); although the optical continuum is very blue, there is a sharp decline in flux density shortward of the Ca II H + K trough. A very noisy CCD spectrum, obtained on Jan. 17, suggests that SN 2004gx (IAUC 8459) is probably of type II. A broad emission feature at the wavelength of Hα is visible in the redshifted spectrum.

H. Navasardyan, S. Benetti, N. Elias-Rosa, and A. Harutunyan, Osservatorio Astronomico di Padova; and A. Pastorello, Max-Planck-Institut für Astrophysik, Garching, on behalf of the European RTN colllaboration (cf. IAUC 7987), report that a spectrum of SN 2005G (cf. IAUC 8465), obtained on Jan. 18.16 UT with the Asiago 1.8-m telescope (+ AFOSC; range 355–780 nm, resolution 2.4 nm), shows it to be a type-Ia supernova ∼ 10 days past maximum. The spectrum closely resembles that of SN 1994D (Patat et al. 1996, MNRAS 278, 111) at comparable phase. The expansion velocity deduced from the Si II 635.5-nm minimum is ∼ 9400 km/s (adopting the NED recession velocity of 6938 km/s for the parent galaxy).

COMET 163P/2004 V4 (NEAT) Following the identification of observations of comet P/2004 V4 (cf.

IAUC 8429, 8438) in 1990–1991 and 1997 (cf. MPC 53257, 53303, 53307), the comet has been numbered 163P.

2005 January 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green